# references.bib

### v4

@@ 1.6 Heisenberg calculus for pseudo-differential operators
    For every $x\in M$ and $\pi\in gr(H)_{x}$, $\sigma(P,x,\pi)$ is injective on $C^{\infty}(H_{\pi})$.
@@
Let $P$ be an H-elliptic differential operator of order $k$. For arbitrary $s\in\mathbb{R}$, $P$ can be extend to the Sobolev space $\tilde{H}^{s}(M)$,
@@
We define the analytical index to be $\text{Ind}(P)=\dim\ker\bar{P}-\dim\ker\bar{P}^{\ast}$. Since H-elliptic operators are hypoelliptic, the kernel of $\bar{P}$ and $\bar{P}^{\ast}$ lie in $C^{\infty}(M)=\bigcap_{s\in\mathbb{R}}\tilde{H}^{s}(M)$, so such analytical index is well-defined.
@@
The analytical index is independent of the choice of $s$ and only depends on the symbol class $[\sigma_{H}(P)]\in K_{0}(C^{\ast}(Gr(H)))$.
Denote the map from the symbol class to the analytical index by
$\text{Ind}_{H^{\bullet}}:K_{0}(C^{\ast}(Gr(H)))\to\mathbb{Z}.$
Corollary 1.6.19.
In the $G$-invariant cases, all the $KK$-classes are replaced by equivariant $KK$-classes, and the analytical index $\text{Ind}(P)=[\text{Ker}\ P]-[\text{Coker}\ P]$ in the representation ring $R(G)$ only depends on the symbol class $[\sigma_{H}(P)]\in K^{G}_{0}(C^{\ast}(Gr(H)))$, i.e
$\text{Ind}_{H^{\bullet}}:K_{0}^{G}(C^{\ast}(Gr(H)))\to R(G).$
@@ 1.7 Groupoid approach and Connes-Thom’s isomorphism
Then, we obtain a commutative diagram
@@
Each component $A_{ij},i,j=0,1$ is a restriction of the global $C^{\ast}$-algebra $A$. The compositions along both side of the commutative diagram from $K_{\ast}(A_{00})$ to $K_{\ast}(A_{11})$ illustrate the induced map on $K$-theory, composition of the lift map from $A_{00}$ to the global algebra $A$, and the restriction map from $A$ to $A_{11}$. The result can then be derived as follows. ∎
@@
The same procedure can be performed by replacing $K_{\ast}(\cdot)$ with $KK_{\ast}(A^{\prime},\cdot)$ for $A^{\prime}$ an arbitrary $C^{\ast}$-algebra.
Theorem 1.7.4 (Connes-Thom’s isomorphism, [nistor2003index]).
@@
There is a $KK$-element $[\mathcal{E}_{0}]\in KK(C^{\ast}(\mathcal{G}),C_{0}(\mathfrak{g}^{\ast}))$, called the Connes-Thom class, such that the $KK$-product with $[\mathcal{E}_{0}]$ over $C^{\ast}(\mathcal{G})$ yields the Connes-Thom’s isomorphism, i.e.
@@
Referencing [fack1981connes], an analog of the standard Connes-Thom’s isomorphism for $C^{\ast}$-dynamic system $(A,\mathbb{R},\alpha)$:
@@
can be represented in the Kasparov product with an element in $KK^{1}(A,A\rtimes_{\alpha}\mathbb{R})$. The isomorphism constructed in [nistor2003index] is a composite of such basic Connes-Thom’s isomorphisms. By associativity, the composition
@@
Let us give a concrete construction of the Connes-Thom’s isomorphism in Theorem 1.7.4. Let $\mathcal{G}$ represent a bundle of nilpotent Lie groups over a compact manifold $M$. We can define a bundle of nilpotent Lie groups over $M\times[0,1]$, denoted by $\mathbb{G}$ and called the adiabatic groupoid for $\mathcal{G}$:
@@
It is also called the Connes-Thom’s isomorphism. We will use this isomorphism in Chapter 3 and 4 in the H-elliptic theory and the transversally H-elliptic theory.
@@ 2.1 Elliptic operators from $KK$-theoretic viewpoint
This section introduced the index theory of elliptic operators discussed in [kasparov2017elliptic], elucidating the symbol class and index class in $KK$-homology.
@@
  • The symbol class for a zero-order elliptic operator is defined by the $KK$-pairing
    $[\sigma(P)]=[(C_{0}(T^{\ast}M;p^{\ast}(E)),\sigma(P)]\in K_{0}^{G}(C_{0}(T^{\ast}M)),$
@@
    $[P]=[(L^{2}(E),P)]\in K_{G}^{0}(\mathbb{C}).$
@@
Let $M$ be a compact Riemannian manifold and $G$ be a Lie group acting isometrically on $M$. Let $P$ be a bounded elliptic operator of order 00 on $M$, then
$[P]=[\sigma(P)]\otimes_{C_{0}(T^{\ast}M)}[\mathcal{D}_{M}]\in KK_{G}(\mathbb{C},\mathbb{C}),$
where $[\mathcal{D}_{M}]\in K^{0}_{G}(C_{0}(TM))$ denotes the $KK$-class of the Dolbeault operator, the $KK$-product on the right-hand side is considered that
$KK_{G}(\mathbb{C},C_{0}(T^{\ast}M))\times KK_{G}(C_{0}(T^{\ast}M),\mathbb{C})\xrightarrow{{\otimes}_{C_{0}(T^{\ast}M)}}KK_{G}(\mathbb{C},\mathbb{C}).$
Here, the Dolbeault operator represents the canonical Dirac operator $D$ on $TM$ with the symbol
$\sigma_{D}(x,\xi,\chi,\zeta)=(\text{ext}\ \zeta+\text{int}\ \zeta)+i(\text{ext}\ \chi-\text{int}\ \chi)).$
@@
The proof is given in [kasparov2022k], and we briefly conclude the proof here.
Primarily, the proof constructs a rotation homotopy between the symbol of two operators, representing the $KK$-classes of the symbol class and index class.
More explicitly, considering the Kasparov product of the Bott–Dirac element $[\mathcal{C}]$ and $[P]$ yields an operator that consolidates $\mathcal{C}\hat{\otimes}1$ and the lifting $\tilde{P}$ of the operator $P$ to $TM$, as $\tilde{P}$ denotes a $KK$-theoretic connection of $P$. The resulting operator acts on $TM$.
The Kasparov product of $\sigma_{F}$ and $[\mathcal{D}_{M}]$ can be represented as the operator combining the lifting $\widetilde{\mathcal{D}_{M}}$ and $1\hat{\otimes}\sigma_{F}$, functioning on $T^{\ast}M$.
Through a rotation of $(\xi,\chi)$-coordinate, the symbols of these operators are rendered homotopic. Given that the Bott–Dirac element $[\mathcal{C}]$ represent the trivial element in $KK$-homology, we see that
@@ 2.2 Transversally elliptic operators from $KK$-theoretic viewpoint
For transversally elliptic operators defined in Definition 2.2.2, it turns out that the direct replacement of the $C_{0}(TM)$ space by $C_{0}(T_{G}M)$ space is not enough. As discussed in [berline1996chern], additional conditions are necessary to identify the “good” symbols due to the inadequate nature of the algebra $C_{0}(T_{G}M)$ for formulating an index theorem.
Kasparov’s index theorem for transversally elliptic operators in $KK$-theory is introduced in [kasparov2022k], presenting his transversally elliptic theory, including the concept of the tangent Clifford symbol, which underpins the definition of transversal H-ellipticity.
Let $M$ be a compact Riemannian manifold equipped with a smooth, isometric action of a Lie group $G$. Let $x\in M$ and $f_{x}:G\to M$ illustrate the map defined by $g\mapsto g(x)$. The Lie algebra of the group $G$ is denoted as $\mathfrak{g}$, and accordingly, $f^{\prime}_{x}:\mathfrak{g}\to T_{x}M$, which accounts for the derivative of $f_{x}$ at $e$. The $G$-action on $M$ is proper, and thus, there exists a $G$-invariant Riemannian metric on $\mathfrak{g}_{M}=M\times\mathfrak{g}$. There is a $G$-invariant quadratic form $q$ on covectors $\xi\in T_{x}^{\ast}M$ by
@@
and a $G$-invariant map
$\varphi_{x}:=f^{\prime}_{x}{f^{\prime}_{x}}^{\ast}:T_{x}M\to T_{x}M.$
@@
A $G$-invariant pseudo-differential operator $P$ of order 00 on a vector bundle $E$ over $M$ is called transversally elliptic if its symbol $\sigma(P)$ satisfies the following condition: for any $\epsilon\geq 0$, there exists $c>0$ such that for any $x\in M$, $\xi\in T^{\ast}_{x}M$,
@@
  • For a $G$-invariant properly supported normalized pseudo-differential operator $P$ of order 00, the original definition of the transversally elliptic is that if $\lVert\sigma(P,x,\xi)^{2}-1\rVert\to 0$ uniformly in $x\in M$ as $(x,\xi)\to\infty$ in $T_{G}M$. Note that we have replaced the operator $P$ by the normalization
@@
    such that it acts on the $\mathbb{Z}_{2}$-graded bundle $E=E_{0}\oplus E_{1}$ of 00-order and self-adjoint. By [kasparov2017elliptic], Lemma 6.2, these two definitions coincide for 00-order pseudo-differential operators.
@@
    for some certain $c>0$. The symbol $\sigma(P)$ of a transversally elliptic operator will induce a class $[\sigma(P)]\in KK_{G}(\mathbb{C},\mathcal{S}_{tr}(M))$.
@@
Let $V$ be a real vector bundle over $M$ equipped with a $G$-invariant Riemannian metric. Take $\text{Cliff}(V,Q)$ the Clifford algebra bundle associated with the quadratic form $Q(v)=\lVert v\rVert$ on $V$. The norm on fibers of the Clifford bundle $\text{Cliff}(V,Q)$ is defined by its embedding into the bundle of endomorphisms $\mathcal{L}(\Lambda^{\ast}(V))$ of the exterior algebra of $V$.
@@
  • Define $Cl_{V}(M)$ to be the complexification of the algebra of continuous sections of $\text{Cliff}(V,Q)$ over $M$, forming a $C^{\ast}$-algebra with the sup-norm on sections.
@@
    The index class $[P]$ is defined as $[(L^{2}(E),P)]\in K^{\ast}(C^{\ast}(G))$.
@@
    $[\sigma^{tcl}(P)]:=[(C_{0}(T^{\ast}M;p^{\ast}E)\hat{\otimes}_{C(M)}Cl_{\Gamma}(M),\sigma^{tcl}(P))]\in KK_{G}(\mathbb{C},Cl_{\Gamma}(TM)),$
@@
    $\sigma^{tcl}(P)(x,\xi):=\sigma(P)(x,\xi)\hat{\otimes}1+(1-\sigma(P)^{2})^{1/2}\hat{\otimes}f_{\Gamma}(x,\xi)\in\mathcal{L}((C_{0}(T^{\ast}M;p^{\ast}E)\hat{\otimes}_{C(M)}Cl_{\Gamma}(M)),$
@@
    • $Cl_{\Gamma}(TM):=Cl_{\Gamma}(M)\otimes_{C(M)}C_{0}(TM)$,
@@
    The tangent-Clifford symbol class is denoted by $[\sigma^{tcl}(P)]\in KK_{G}(\mathbb{C},Cl_{\Gamma}(TM))$.
@@
By the $KK^{G}$-equivalence in some sense between the algebras $Cl_{\Gamma}(TM)$ and $\mathcal{S}_{tr}(M)$([kasparov2022k]), a one-to-one correspondence exists between the symbol class $[\sigma(P)]\in KK_{G}(\mathbb{C},\mathcal{S}_{tr}(M))$ and the tangent-Clifford symbol class, $[\sigma^{tcl}(P)]\in KK_{G}(\mathbb{C},Cl_{\Gamma}(TM))$.
The definition of $\mathcal{S}_{tr}(M)$ space depends on the inequality estimate, rendering it unsuitable for the transversally H-elliptic case as $C^{\ast}(Gr(H))$ is noncommutative. However, the tangent-Clifford symbol class only depends on the orbit direction in $T_{x}M$ for each $x\in M$, which is considerable in the transversally H-elliptic case.
Based on the $KK^{G}$-equivalence, we can always analyze the transversally elliptic symbols by using the tangent-Clifford symbol class $[\sigma^{tcl}(P)]$ in the following contents.
@@
The Clifford Dolbeault element $[D^{cl}_{M,\Gamma}]\in K^{0}(C^{\ast}(G,Cl_{\Gamma}(TM)))$ is defined to be the $KK$-element $[(\mathcal{H},\pi,F)]$, where
@@
  • $F=\mathcal{D}_{M}$ is the same operator as the Dolbeault element $[\mathcal{D}_{M}]\in K^{0}_{G}(C_{0}(TM))$;
@@
Let $M$ be a compact Riemannian manifold and $G$ be a Lie group which acts on $M$ isometrically, Let $P$ represent a $G$-invariant, $L^{2}$-bounded transversally elliptic operator on $M$ of order 00, then
@@
where $j^{G}$ denote the natural map from $G$-equivariant $KK$-homology to $KK$-homology of crossed product with $G$, refer to [blackadar1998k]. The $KK$-product on the right-hand side is considered
$KK(C^{\ast}(G),C^{\ast}(G,Cl_{\Gamma}(TM))\times KK(C^{\ast}(G,Cl_{\Gamma}(TM),\mathbb{C})\xrightarrow{\otimes_{C^{\ast}(G,Cl_{\Gamma}(TM)}}KK(C^{\ast}(G),\mathbb{C}).$
@@
In this chapter, we introduce the H-elliptic index theorem. In [mohsen2022index], Mohsen constructed a diagram of groupoids over $[0,1]^{2}$, such that the groupoid structure on each side independently forms a Lie groupoid structure. A $C^{\ast}$-algebra is defined on this groupoid, analyzed and integrated to deduce the index theorem for H-elliptic differential operators.

### v5

@@ 1.7 Index map for H-elliptic operators
    For every $x\in M$ and $\pi\in gr(H)_{x}$, $\sigma_{H}(P,x,\pi)$ is injective on $C^{\infty}(H_{\pi})$.
@@
Let $P$ be an H-elliptic differential operator of order $k$. For arbitrary $s\in\mathbb{R}$, $P$ can be extended to the Sobolev space $\tilde{H}^{s}(M)$,
@@
We define the analytical index of $P$ to be $\text{Ind}(P)=\dim\ker\bar{P}-\dim\ker\bar{P}^{\ast}$.
Since H-elliptic operators are hypoelliptic, the kernel of $\bar{P}$ and $\bar{P}^{\ast}$ lie in $C^{\infty}(M)=\bigcap_{s\in\mathbb{R}}\tilde{H}^{s}(M)$, so the analytical index is well-defined.
@@
Let $M$ be a compact filtered manifold. The analytical index is independent of the choice of $s$ and only depends on the symbol class $[\sigma_{H}(P)]\in KK(C(M),C^{\ast}(Gr(H)))$. The map from the symbol class to the analytical index is denoted by the composition
$\text{Ind}_{H^{\bullet}}:KK(C(M),C^{\ast}(Gr(H)))\to K^{0}(C(M))\to\mathbb{Z},$
where the second map is induced from the inclusion $\mathbb{C}\subset C(M)$.
In the $G$-invariant case, all the $KK$-classes are replaced by equivariant $KK$-classes, and the analytical index $\text{Ind}(P)=[\text{Ker}\ P]-[\text{Coker}\ P]$ in the representation ring $R(G):=KK^{G}(\mathbb{C},\mathbb{C})$ only depends on the symbol class $[\sigma_{H}(P)]\in KK^{G}(C(M),C^{\ast}(Gr(H)))$, i.e, we have the analytical index map
$\text{Ind}_{H^{\bullet}}:KK^{G}(C(M),C^{\ast}(Gr(H)))\to K^{0}(C^{\ast}(G,C(M)))\to R(G).$
@@ 1.8 Groupoid approach and Connes-Thom’s isomorphism
In this section, we explain the groupoid approach which was invented by Connes in [connes1994noncommutative]. We will also introduce the Connes-Thom’s isomorphism, which is important in van Erp and Mohsen’s index theory for H-elliptic operators.
@@
Then, we obtain a commutative diagram:
@@
Each component $A_{ij},i,j=0,1$ is a restriction of the global $C^{\ast}$-algebra $A$. The compositions along both side of the commutative diagram from $K_{\ast}(A_{00})$ to $K_{\ast}(A_{11})$ illustrate the induced map on $K$-theory, composition of the lift map from $A_{00}$ to the global algebra $A$, and the restriction map from $A$ to $A_{11}$. The result then follows. ∎
@@
The same procedure can be performed by replacing $K_{\ast}(\cdot)$ with $KK(A^{\prime},\cdot)$ for $A^{\prime}$ an arbitrary $C^{\ast}$-algebra.
Theorem 1.8.4 (Connes-Thom’s isomorphism, [nistor2003index], Lemma 4.3).
@@
This Connes-Thom’s isomorphism can be written as the $KK$-product with some special element as follows:
@@
There is a $KK$-element $[\mathcal{E}_{0}]\in KK(C^{\ast}(\mathcal{G}),C_{0}(\mathfrak{g}^{\ast}))$, called the Connes-Thom class, such that the $KK$-product with $[\mathcal{E}_{0}]$ over $C^{\ast}(\mathcal{G})$ yields the Connes-Thom’s isomorphism in Theorem 1.8.4, i.e.
@@
By Proposition 1 in [fack1981connes], the standard Connes-Thom’s isomorphism for $C^{\ast}$-dynamic system $(A,\mathbb{R},\alpha)$:
@@
can be represented as the Kasparov product with an element in $KK^{1}(A,A\rtimes_{\alpha}\mathbb{R})$. By the construction in the proof of Lemma 4.3 in [nistor2003index], the Connes-Thom’s isomorphism in Theorem 1.8.4 is a composite of such standard Connes-Thom’s isomorphisms. By the associativity of $KK$-products, we see that the composition
@@
Now we give a concrete construction for the Connes-Thom’s isomorphism in Theorem 1.8.4. Let $\mathcal{G}$ represent a bundle of nilpotent Lie groups over a locally compact manifold $M$. We can define a bundle of nilpotent Lie groups over $M\times[0,1]$, denoted by $\mathbb{G}$ and called the adiabatic groupoid for $\mathcal{G}$:
@@
It is also called the Connes-Thom’s isomorphism. It will be used in Chapter 3 in Mohsen’s H-elliptic index theory.
@@
In order to give an appropriate definition for transversally H-elliptic operators from H-elliptic theory, we first observe the differences between the elliptic and transversally elliptic operators in $KK$-theory. We refer the elliptic theory to [kasparov2017elliptic] and the transversal elliptic theory to [kasparov2022k], so that we can avoid the errors in transversal elliptic theory in [kasparov2017elliptic], which has been clarified by Kasparov in [kasparov2022k]. We will focus on the index of operators in $KK_{0}$, the $KK_{1}$ index for self-adjoint operators on non-graded bundles can be analogously analyzed.
@@ 2.1 Elliptic operators from $KK$-theoretic viewpoint
In this section, we introduce the index theory of elliptic operators discussed in [kasparov2017elliptic], elucidating the symbol class and index class in $KK$-homology.
@@
  • The symbol class for a 00-order $G$-invariant elliptic operator is defined by the $KK$-pairing
    $[\sigma(P)]=[(C_{0}(T^{\ast}M;p^{\ast}(E)),\sigma(P)]\in KK^{G}(C_{0}(M),C_{0}(T^{\ast}M)),$
@@
    $[P]=[(L^{2}(E),P)]\in K_{G}^{0}(C_{0}(M)).$
By estimate (*2.1.2), these definitions are well-defined.
@@
Let $M$ be a complete Riemannian manifold and $G$ be a locally compact Lie group acting properly and isometrically on $M$. Let $P$ be a $G$-invariant bounded elliptic operator of order 00 on $M$, then
$[P]=[\sigma(P)]\otimes_{C_{0}(T^{\ast}M)}[\mathcal{D}_{M}]\in K^{0}_{G}(C_{0}(M)),$
where $[\mathcal{D}_{M}]\in K^{0}_{G}(C_{0}(TM))$ denotes the $KK$-class of the Dolbeault operator, the $KK$-product on the right-hand side is considered over
$KK^{G}(C_{0}(M),C_{0}(T^{\ast}M))\times KK^{G}(C_{0}(T^{\ast}M),\mathbb{C})\xrightarrow{{\otimes}_{C_{0}(T^{\ast}M)}}K^{0}_{G}(C_{0}(M)).$
Here, the Dolbeault operator represents the canonical Dirac operator $D_{M}$ on $TM$ with the symbol
$\sigma_{D_{M}}(x,\xi,\chi,\zeta)=(\text{ext}\ \zeta+\text{int}\ \zeta)+i(\text{ext}\ \chi-\text{int}\ \chi)).$
@@
The proof is given in [kasparov2022k], and we briefly conclude it here.
It is important to note that the operator $P$ is differential but the symbol $\sigma(P)$ is a multiplication operator. So we shall construct a rotation homotopy relating the operator and symbol.
More explicitly, the Kasparov product of the Bott–Dirac element $[\mathcal{C}]$ and $[P]$ yields an operator on $TM$ that combines $\mathcal{C}\hat{\otimes}1$ and the lifting $\tilde{P}$ of the operator $P$ to $TM$, where $\tilde{P}$ denotes a $KK$-theoretic connection of $P$. The resulting operator acts on $TM$. The Kasparov product of $[\sigma_{F}]$ and $[\mathcal{D}_{M}]$ can be represented as the operator combining the lifting $\widetilde{\mathcal{D}_{M}}$ and $1\hat{\otimes}\sigma_{F}$, functioning on $T^{\ast}M$.
The symbols of these two operators are homotopic through a rotation homotopy of $(\xi,\chi)$-coordinate. The Bott–Dirac element $[\mathcal{C}]$ represent the trivial element in $KK$-homology, so we see that
@@ 2.2 Transversally elliptic operators from $KK$-theoretic viewpoint
For transversally elliptic operators, it turns out that the direct replacement of the $C_{0}(TM)$ space by $C_{0}(T_{G}M)$ space is not enough. As discussed in [berline1996chern], additional conditions are necessary to consider the “good” symbols due to the inadequate nature of the algebra $C_{0}(T_{G}M)$ for formulating an index theorem.
Kasparov’s index theorem for transversally elliptic operators in $KK$-theory is introduced in [kasparov2022k], including the concept of the tangent-Clifford symbol, which underpins our definition of transversal H-ellipticity.
Let $M$ be a complete Riemannian manifold equipped with a smooth, proper, and isometric action of a locally compact Lie group $G$. Let $x\in M$ and $f_{x}:G\to M$ illustrate the map defined by $g\mapsto g(x)$. The Lie algebra of the group $G$ is denoted as $\mathfrak{g}$, and accordingly, $f^{\prime}_{x}:\mathfrak{g}\to T_{x}M$, which accounts for the derivative of $f_{x}$ at $e$. The $G$-action on $M$ is proper, and thus, there exists a $G$-invariant Riemannian metric on $\mathfrak{g}_{M}=M\times\mathfrak{g}$. There is a $G$-invariant quadratic form $q$ on covectors $\xi\in T_{x}^{\ast}M$ by
@@
and we have the following $G$-invariant map
$\varphi_{x}:=(f^{\prime}_{x}{f^{\prime}_{x}}^{\ast})^{1/2}:T_{x}M\to T_{x}M.$
@@
A $G$-invariant pseudo-differential operator $P$ of order 00 on a vector bundle $E$ over $M$ is called transversally elliptic if its symbol $\sigma(P)$ satisfies the following condition: for any $\epsilon>0$, there exists $c>0$ such that for any $x\in M$, $\xi\in T^{\ast}_{x}M$,
@@
  • For a $G$-invariant properly supported normalized pseudo-differential operator $P$ of order 00, the original definition of the transversally elliptic is that $\lVert\sigma(P,x,\xi)^{2}-1\rVert\to 0$ uniformly in $x\in M$ as $(x,\xi)\to\infty$ in $T_{G}M$. Note that we have replaced the operator $P$ by the normalization
@@
    such that it acts on the $\mathbb{Z}_{2}$-graded bundle $E=E_{0}\oplus E_{1}$ of 00-order and is self-adjoint. By [kasparov2017elliptic], Lemma 6.2, these two definitions coincide for 00-order pseudo-differential operators.
@@
    for some certain $c>0$. The symbol $\sigma(P)$ of a transversally elliptic operator will induce a class $[\sigma(P)]\in KK^{G}(C_{0}(M),\mathcal{S}_{tr}(M))$.
@@
Let $V$ be a real vector bundle over $M$ equipped with a $G$-invariant Riemannian metric. Take $\text{Cliff}(V,Q)$ the Clifford algebra bundle associated with the quadratic form $Q(v)=\lVert v\rVert^{2}$ on $V$. The norm on fibers of the Clifford bundle $\text{Cliff}(V,Q)$ is defined by its embedding into the bundle of endomorphisms $\mathcal{L}(\Lambda^{\ast}(V))$ of the exterior algebra of $V$.
@@
  • Define $Cl_{V}(M)$ to be the complexification of the algebra of continuous sections of $\text{Cliff}(V,Q)$ over $M$ vanishing at infinity of $M$, forming a $C^{\ast}$-algebra with the sup-norm on sections. We will use a particular case of this algebra when $V=\tau=T^{\ast}M$. This algebra will be denoted $Cl_{\tau}(M)$.
@@
Similar to Definition 2.1.3, we can define the index class and symbol class for transversally elliptic operators.
@@
    The index class $[P]$ is defined as $[(L^{2}(E),P)]\in K^{\ast}(C^{\ast}(G,C_{0}(M)))$.
@@
    $[\sigma^{tcl}(P)]:=[(C_{0}(T^{\ast}M;p^{\ast}E)\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M),\sigma^{tcl}(P))]\in KK^{G}(C_{0}(M),Cl_{\Gamma}(TM)),$
@@
    $\sigma^{tcl}(P)(x,\xi):=\sigma(P)(x,\xi)\hat{\otimes}1+(1-\sigma(P)^{2})^{1/2}\hat{\otimes}f_{\Gamma}(x,\xi)\in\mathcal{L}((C_{0}(T^{\ast}M;p^{\ast}E)\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M)),$
@@
    • $Cl_{\Gamma}(TM):=Cl_{\Gamma}(M)\otimes_{C_{0}(M)}C_{0}(TM)$,
@@
    The tangent-Clifford symbol class is denoted by $[\sigma^{tcl}(P)]\in KK^{G}(C_{0}(M),Cl_{\Gamma}(TM))$.
@@
By the $KK^{G}$-equivalence in some sense between the algebras $Cl_{\Gamma}(TM)$ and $\mathcal{S}_{tr}(M)$ (see Theorem 7.10 in [kasparov2022k]), there is a one-to-one correspondence between the symbol class $[\sigma(P)]\in KK^{G}(C_{0}(M),\mathcal{S}_{tr}(M))$ and the tangent-Clifford symbol class, $[\sigma^{tcl}(P)]\in KK^{G}(C_{0}(M),Cl_{\Gamma}(TM))$. We will always analyze the transversally elliptic symbols by using the tangent-Clifford symbol class $[\sigma^{tcl}(P)]$ in the following contents.
The definition of $\mathcal{S}_{tr}(M)$ space depends on the inequality estimate, rendering it unsuitable for the transversally H-elliptic case as $C^{\ast}(Gr(H))$ is noncommutative. On the other hand, the tangent-Clifford symbol class only depends on the orbit direction in $T_{x}M$ for each $x\in M$, which is considerable in the transversally H-elliptic case, and we will use it to give the definition in Section 4.5.
@@
The Clifford-Dolbeault element $[D^{cl}_{M,\Gamma}]\in K^{0}(C^{\ast}(G,Cl_{\Gamma}(TM)))$ is defined to be the $KK$-element $[(\mathcal{H},\pi,F)]$, where
@@
  • $F=\mathcal{D}_{M}$ is the same operator as the Dolbeault element $[\mathcal{D}_{M}]\in K^{0}_{G}(C_{0}(TM))$, which has been defined in Theorem 2.1.4;
@@
Let $M$ be a complete Riemannian manifold and $G$ be a locally Lie group which acts on $M$ properly and isometrically, Let $P$ represent a $G$-invariant, $L^{2}$-bounded transversally elliptic operator on $M$ of order 00, then
@@
where $j^{G}$ denote the natural map from $G$-equivariant $KK$-homology to $KK$-homology of crossed product with $G$, refer to [blackadar1998k]. The $KK$-product on the right-hand side is considered over
$KK(C^{\ast}(G,C_{0}(M)),C^{\ast}(G,Cl_{\Gamma}(TM))\times K^{0}(C^{\ast}(G,Cl_{\Gamma}(TM))\xrightarrow{\otimes_{C^{\ast}(G,Cl_{\Gamma}(TM)}}K^{0}(C^{\ast}(G,C_{0}(M))).$
@@
In Chapter 2, we have seen the difference between elliptic and transversally elliptic operators. Before defining the transversally H-elliptic operators, we introduce the H-elliptic index theory and write the index theorem from the $KK$-theoretic viewpoint.